OpenStudy (anonymous):

Solve using the square root method. Round to the nearest tenth if necessary. x^2 + 15 =0 A. {3.9} B. {3.9, -3.9} C. {7.5, -7.5} D. no real solution

jimthompson5910 (jim_thompson5910):

\[\Large x^2 + 15 = 0\] \[\Large x^2 + 15-15 = 0-15\] \[\Large x^2 = -15\] \[\Large \sqrt{x^2} = \sqrt{-15}\] \[\Large x = \sqrt{-15}\] So there are no real solutions.

OpenStudy (mertsj):

Subtract 15 from both sides and then take the square root of both sides.